three no. are in the ratio 3:4:5 if the sum of the largest and the smallest equal the sum of the third and 52, find the numbers.
Answer by edjones(8007)

sum of the largest and the smallest equal the sum of the third and 52, find the numbers.
the new triangle must have the sides proportional to 3:4:5
let x= the multiple of each side to make a side of the new triangle.
3x+5x=4x+52
3x+5x-4x=52
4x=52
divide 4 into each side: x=13
13*3=39
13*4=52
13*5=65
Check:
39+65=52+52
104=104
Ed